OpenStudy (anonymous):

There are 40 questions on Allison's English test. The multiple choice questions are worth 2 points each, and the short answer questions are worth 5 points each. The entire test is worth 125 points. Let x = multiple choice questions; and let y = short answer questions. How many short answer questions are there on the test? @hartnn @charlotte123

OpenStudy (charlotte123):

2x + 5y = 125

OpenStudy (anonymous):

So do I plug in stuff?

OpenStudy (anonymous):

Never mind. Dunno how to solve this lol..

OpenStudy (charlotte123):

Yep Remember - There are 40 problems on the test :)

OpenStudy (anonymous):

So .. I don't know where to go from here =/

OpenStudy (anonymous):

@charlotte123 What do I do from that equation you gave me??

OpenStudy (charlotte123):

You solve for y :)

OpenStudy (anonymous):

I do this how? =/

OpenStudy (charlotte123):

LOL You really are stuck!

OpenStudy (anonymous):

I am xD I stink at math =/

OpenStudy (charlotte123):

Ah It happens LOL

OpenStudy (charlotte123):

40 problems - 125 POINTS <---- 2x + 5y

OpenStudy (charlotte123):

^ Where 5y represents short answer problems and 2x represents long answer problems.

OpenStudy (anonymous):

Okay. What do I do with the 40? and the 125? O_o

OpenStudy (charlotte123):

LOL

OpenStudy (anonymous):

Cx I hate math gahhh

OpenStudy (charlotte123):

2 POINTS * LONG PROBLEMS + 5 POINTS * SHORT PROBLEMS = 125 POINTS [40 PROBLEMS] <-----

OpenStudy (charlotte123):

I know I know BUT YOU CAN DO IT!

OpenStudy (anonymous):

Ok. Why is the 40 in brackets? Jeesh feel so dumb Cx

OpenStudy (charlotte123):

You are not - At least you try! Some people on here never ask for help and want the answers To the problems Instead - ASKING QUESTIONS MAKES YOU SMART!

OpenStudy (charlotte123):

You see - The problems cannot exceed 40 Means that the problems must add To 125 POINTS WITHOUT EXCEEDING 40 :)

OpenStudy (charlotte123):

I shall hand you a clue - 2x ----> x = 10 there.

OpenStudy (anonymous):

So -20?

OpenStudy (charlotte123):

LOL NO

OpenStudy (anonymous):

Oh wait that's not a negative sign xD 20 oops.

OpenStudy (charlotte123):

10 + -20 ≠ 125

OpenStudy (charlotte123):

OH WAIT I MADE A MISTAKE Forgot about that there are 40 problems - Would have been 33 If I handled It my way LOL

OpenStudy (anonymous):

Ok so that resolves the problem to currently look like, what? ;o

OpenStudy (charlotte123):

GOT IT 2x <---- Clue - x = 25.

OpenStudy (charlotte123):

So - 25 + 5x = 125 <----

OpenStudy (charlotte123):

And 25 + 5x = 40 <---- x Is the same In both parts on these.

OpenStudy (anonymous):

What? xD Here's what I was thinking but is incorrect according to what you're doing Cx. 2(25) = 50 But in your equation 25 + 5(25) = 125 25 + 125 ≠ 125 LOL.. I'm sooooo confuseddddddddd lol!

OpenStudy (charlotte123):

L O L

OpenStudy (charlotte123):

x = 25 25 * 2 = 50 125 - 50 = 75 75/5 = x

OpenStudy (charlotte123):

The 1st one was for the POINTS 2nd one was for the PROBLEMS 25 + x = 40.

OpenStudy (charlotte123):

OH WAIT NOW I SEE

OpenStudy (charlotte123):

-Face Plop-

OpenStudy (charlotte123):

THERE MUST BE CHOICES

OpenStudy (anonymous):

There are no choices =/.. It's "Type your response".

OpenStudy (charlotte123):

WAIT I WAS RIGHT - I JUST MADE A MISTAKE

OpenStudy (charlotte123):

x + y = 40 And 2x + 5y = 125

OpenStudy (charlotte123):

NOT 2x + 5y = 40

OpenStudy (anonymous):

Ok

OpenStudy (charlotte123):

Now - 2x <---- x = 25 here.

OpenStudy (anonymous):

So 50? Cause 2 * 25 = 50?

OpenStudy (charlotte123):

YES So 50 + 5y = 125 And 25 + y = 40

OpenStudy (anonymous):

y = 15?

OpenStudy (charlotte123):

YESSSSSSSSSSSSSSSSSSSSSS

OpenStudy (charlotte123):

THERE YOU GO

OpenStudy (anonymous):

THANK THE MIGHTY GOD OF MATH! (You) Cx

OpenStudy (charlotte123):

15 + 25 = 40 2(25) + 5(15) = 125

OpenStudy (charlotte123):

LOLLLLLLLLLLLLLLL

OpenStudy (anonymous):

Wish I could medal you 1284312931293192 times Cx. I'm gonna write you a testimonial (:. Thank you so much!!!

OpenStudy (charlotte123):

PRAISE THE APPRENTICE

OpenStudy (charlotte123):

OMG NO PROBLEM! -Feels so honored- T_T